OpenStudy (anonymous):

8-19). Simplify each term, if necessary, and then use the distributive property to combine terms, if possible. sqrt(50)+ 3 sqrt(2)

OpenStudy (anonymous):

\[\sqrt{50}=\sqrt{25*2}=\sqrt{25}*\sqrt{2}=5\sqrt{2}\]

OpenStudy (anonymous):

now add treating the sqrt(2) 's like they are variables 5x + 3x = 8x so the answer is \[8\sqrt{2}\]